bdrammeh:

a wire is carrying a 2.45 A current. at what distance from the wire is the magnetic field 1.00 x10^-6 T ?= m

TheSmartOne:

\(\text{B} =\frac{\mu _{0}\text{I}}{2\pi \text{r}}\) B is our magnetic field \(\mu_{0} =4π10^{−7}~\frac{\text{T}\cdot \text{ m}}{\text{A}}\) I is current and r is the shortest distance to the wire

TheSmartOne:

Plug and chug

bdrammeh:

can u plug it in for me

TheSmartOne:

How difficult is it? What part of the plugging in confuses you? Which variable are you confused about?

bdrammeh:

everything

bdrammeh:

nvm

TheSmartOne:

What is the magnetic field that they gave you? It's in the question. What is the current that they give?

TheSmartOne:

I gave you the value of \(\mu_{0}\) a few posts about And so once you put everything into the formula, you just have to multiply and divide and solve for 'r'

bdrammeh:

help me with this one a 2.99 x 10^-6 C charge is moving perpendicular (90) to the earth magnetic field (5.00 x 10^-5 T). If the force on it is 2.14 x 10^-8 N, how far is it moving ? m/s

TheSmartOne:

Let's make a new post for this

bdrammeh:

ok